OpenStudy (brinazarski):

Simplify the expression. (Do not expand the expression in the denominator.) (e^2x − e^−2x)^2 − (e^2x + e^−2x)^2 (e^2x + e^−2x)^2

OpenStudy (brinazarski):

\[\left( e ^{2x}-e ^{-2x} \right)^{2}-\left( e ^{2x}+e ^{-2x} \right)^{2}\] ________________ \[\left( e ^{2x}+e ^{-2x} \right)^{2}\]

OpenStudy (brinazarski):

Just to make it clearer

OpenStudy (anonymous):

``` \[ \frac{(e^{2x}-e^{-2x})^2-(e^{2x}+e^{-2x})^2}{(e^{2x}+e^{-2x})^2} \] ``` \[ \frac{(e^{2x}-e^{-2x})^2-(e^{2x}+e^{-2x})^2}{(e^{2x}+e^{-2x})^2} \]

OpenStudy (brinazarski):

I figured it out just now! ^^;

OpenStudy (anonymous):

Let \(a=e^{2x}-e^{-2x}\) and \(b=e^{2x}+e^{-2x}\)\[ \frac{a^2-b^2}{b^2} \]

OpenStudy (anonymous):

This is difference of squares.\[ \frac{a^2-b^2}{b^2}=\frac{(a-b)(a+b)}{b^2} \]

OpenStudy (brinazarski):

I understand it now ^^;